OpenStudy (anonymous):

simplify x/7x+x^2 answer choices: 1/7x+x; where x≠-7 1/7x; where x≠0 1/7+x; where x≠0, -7 1/7

Nnesha (nnesha):

factor denominator

OpenStudy (anonymous):

btw these are all fraction

OpenStudy (anonymous):

and i know how to simplify it, but why does it have the: where x≠??

Nnesha (nnesha):

alright what did you get at the denominator 7x+x^2 = ?

OpenStudy (anonymous):

x(x+7) ?

OpenStudy (anonymous):

wait, x+7?

Nnesha (nnesha):

yep right now set it like this \[\huge\rm x(x+7) \cancel{=}0\] denominator can't be equal to 0 bec when you can't divide anything by 0 x/0 = undefined

Nnesha (nnesha):

\[\huge\rm x(x+7) \cancel{=}0\] solve for x

Nnesha (nnesha):

set both x and (x+7)not equal to 0 \[x \cancel{=} 0 , (x+7) \cancel{=} 0\] x can't be equal to 0 \[ \ (x+7) \cancel{=} 0\] solve for x^

OpenStudy (anonymous):

so when it says x≠, i just have to put the numbers that come after is in for x in the equation and see if it works itself out or not?

Nnesha (nnesha):

well x≠something for example lest's say x≠ 6 that's mean xcan't be equal to 6 bec x- 6 when u plug in 6 for x you will get 0 6-6 = 0 that's what we don't want DENOMINATOR can't be equal to 0

OpenStudy (anonymous):

ohhhhhhh okay... thank you soooo much!

Nnesha (nnesha):

np :-)